Find the length of the side of a square if the diagonal is 3 ft. longer than a side.
Answer by nerdybill(5404)

Find the length of the side of a square if the diagonal is 3 ft. longer than a side.
.
Let x = length of the side of a square
then from Pythagorean theorem we get:
x^2 + x^2 = x+3
2x^2 = x+3
2x^2-x-3 = 0
(2x-3)(x+1) = 0
.
x = {3/2, -1}
We can throw out the negative solution leaving:
x = 3/2 feet (or 1 and a half feet)